a. Obtain the linear trend equation for the following data on new checking accounts at Fair Savings Bank and use it to predict expected new checking accounts for periods 16 through 19. (Round your intermediate calculations and final answers to 2 decimal places.)

Period New Accounts Period New Accounts Period New Accounts
1 200 6 237 11 281
2 212 7 249 12 275
3 211 8 250 13 287
4 222 9 255 14 288
5 235 10 267 15 313


Y = + t
Y16 =
Y17 =
Y18 =
Y19 =


b.
Use trend-adjusted smoothing with α = .2 and β = .1 to smooth the new account data in part a. What is the forecast for period 16? Compute the initial trend estimate (Tt) for Period 5 as follows: (Period 4 data – Period 1 data) / 3. Then compute the initial trend-adjusted forecast (TAFt) for Period 5 as follows: Period 4 data + Initial trend estimate for Period 5. Then compute all remaining values (including the St value for Period 5) using the textbook formulas or Excel template. (Round the "Trend"values (Tt) to 3 decimal places and all other intermediate forecast values (TAFt and St) to 2 decimal places. Round your final answer to 2 decimal places.)


Forecast for period 16_____   
Answer #1

Dear student, we cannot able to post solution more than one question as per our policy.

Period (t) New Accounts (Y) t2 tY
1 200 1 200
2 212 4 424
3 211 9 633
4 222 16 888
5 235 25 1175
6 237 36 1422
7 249 49 1743
8 250 64 2000
9 255 81 2295
10 267 100 2670
11 281 121 3091
12 275 144 3300
13 287 169 3731
14 288 196 4032
15 313 225 4695
Total 120 3782 1240 32299
Sign Σt ΣY Σt2   ΣtY

b = [N (ty) – Et EY] /[N (t) - (Et) 2] b = [15*32299 - 120*3782] /[15*1240 - 120°) b = [484485-453840] / [18600 - 14400] b= 3